OpenStudy (anonymous):

A rectangle that is 8 inches long and 5 inches wide has the same area as a triangle that is 2 inches high. What is the base of the triangle? Show and explain each step to find the answer.

OpenStudy (anonymous):

 So the base of the triangle would be 40 ,, 40*2/2 = 40 (the area) which is the same area as the recangle ( 8*5=40 ) ?